# No Compromise in Solution Quality: Speeding Up Belief-dependent Continuous POMDPs via Adaptive Multilevel Simplification

### v1

@@ 7.2 Entropy Estimators and Bounds Study
In this section, we experiment with a passive case of the continuous 2D Light Dark problem from Section 7.1.1. Our goal is to study the various entropy estimators and our derived bounds from Section 6 over the estimator developed in Boers et al. (2010). In this study, we manually supply the robot with an action sequence to conduct. This results in a single lace of the beliefs corresponding to observations that the robot actually obtained by executing a given externally action sequence.
@@
The results for the first and second scenarios are provided in Figs. 9 and 10, respectively. For both scenarios we observe that the bounds become tighter as the number of particles of simplified belief $n^{s}_{x}$ increases. We also witness that all estimators vary but the overall trend is similar, putting aside the discrete entropy over the weights. The discrete entropy over the weights fails to adequately represent the uncertainty of the belief. This is an anticipated result. Let us proceed to the planning experiments.
@@ 7.3.1 Acceleration measures
$\frac{\sum_{i}\Big{(}n_{i,x}^{2}-n_{i,x}^{s}n_{i,x}\Big{)}}{\sum_{i}n_{i,x}^{2}}\cdot 100=\frac{\sum_{i}\Big{(}n_{i,x}-n_{i,x}^{s}\Big{)}}{\sum_{i}n_{i,x}}\cdot 100,$ (61)
where the summation is over the future posterior beliefs in all the belief trees in a number of a consecutive planning sessions in particular scenario. Eq. (61) measures relative speedup without time spent on resimplifications. It is calculated at the end of several consecutive planning sessions. To calculate speedup according to (61) one shall pick up the final number of particles of simplified belief $n^{s}_{i,x}$ used for the simplified reward for each belief node $i$, sum over all the nodes of the belief trees (given or constructed on the fly) from planning sessions, make a calculation portrayed by (61). Importantly, acceleration measure (61) assumes that time of evaluating the motion and observation models do not vary from one evaluation to another.
@@
If the quantities (61) and (62) are identical we can conclude that there will be no overhead from resimplifications and adapting the bounds. Note also that in the first place it is not clear how many particles $n_{x}$ for belief representation to take. The number of particles $n_{x}$ shall be as large as possible due to fact that we do not know when the belief represented by weighted particles will converge to the corresponding theoretical belief. To rephrase that, we always can make the overhead from resimplifications and adaptations neglectable by increasing $n_{x}$.
@@ 7.3.3 Results for 2D Continuous Target Tracking in the Setting of a Given Belief Tree
Our action space is $\mathcal{A}=\{\rightarrow,\nearrow,\uparrow,\nwarrow,\leftarrow,\swarrow,\downarrow,\searrow,Null\}$, where action $Null$ means that agent doesn’t take any action. In this problem we selected the parameters to be $d_{\text{min}}=0.0001$, $\Sigma_{T}=I\cdot\sigma_{T}$ where $\sigma_{T}=0.1$ and $\sigma_{\mathcal{O}}=0.1$, $\gamma=0.95$.
@@ 7.3.4 Experiments with MCTS
In an anytime setting of MCTS we focus on the $2D$-continuous light dark problem from Section 7.1.1. We place a single “light beacon” in the continuous world. Here we changed the reward. The agent’s goal is to get to location $(0,0)$ and execute the terminal action - $Null$. Executing it within a radius of $0.5$ from $(0,0)$ will give the agent a reward of $200$, and executing it outside the radius will yield a negative reward of $-200$. For all other actions the multi-objective reward function is $\rho(b,a,z,b^{\prime})=-\underset{x\sim b^{\prime}}{\mathbb{E}}[\|x\|_{2}]-\hat{\mathcal{H}}(b,a,z,b^{\prime})$. The agent can move in eight evenly spread directions $\mathcal{A}=\{\rightarrow,\nearrow,\uparrow,\nwarrow,\leftarrow,\swarrow,\downarrow,\searrow,Null\}$. Motion and observation models, and the initial belief are $\mathbb{P}_{T}(\cdot|x_{,}a)=\mathcal{N}(x+a,\Sigma_{T})$, $\mathbb{P}_{Z}(z|x)=\mathcal{N}(x,\min\{1,\parallel x-x^{b}\parallel^{2}_{2}\}\cdot\Sigma_{\mathcal{O}})$, $b_{0}=\mathcal{N}(x_{0},\Sigma_{0})$ respectively. $x^{b}$ is the 2D location of the beacon and all covariance matrices are diagonal (i.e. $\Sigma=I\cdot\sigma^{2}$).
@@
We selected the following parameters $x_{0}=\begin{pmatrix}-5.5\\ 0.0\end{pmatrix},\Sigma_{0}=\begin{pmatrix}0.2&0.0\\ 0.0&0.2\end{pmatrix},\sigma_{T}=\sigma_{\mathcal{O}}=0.075$. We experiment with $10$ different configurations (rows of Table 7) that differ in $n_{x}$ (number of particles), $L$ (MCTS simulation depth), and ##\##iter (number of MCTS simulation iterations per planning session). Each scenario comprises $10$ planning sessions, i.e. the agent performs up to $10$ planning action-executing iterations. The scenario stops if the best action determined in planning is $Null$. We repeat each experiment 25 times. In each such repetition we run PFT-DPW and SITH-PFT with the same seed and calculate the relative time speedup in percentage according to (62) where $t_{\mathrm{PFT-DPW}}$ and $t_{\mathrm{SITH-PFT}}$ are running times of a baseline and our methods respectively.
In all different configurations, we obtained significant time speedup of approximately $20\%$ while achieving the exact same solution compared to PFT. In Table 7 we report the mean and standard error of (62) as well as maximum and minimum value. Remarkably, we observe that we never slowdown the PFT-DPW with SITH-PFT. We also present total running times of $25$ repetitions of at most $10$ (the simulation stops if best identified action is $Null$) planning sessions in Table 8. Note that we divided the total planning time by the number of planning sessions in each repetition.
@@ 7.4 Discussion
Although the speedup was significant and steady for all simulations, we did not observe growth in speed-up with growth of number of belief particles in any simulation. This can be explained by the fact that the estimators are far from convergence and increasing number of particles just randomly changes the estimator itself due to sampling induced variance. The limitation of our approach is that it leans on converging bounds, which are not trivial to derive and specific for a particular reward function. In addition, it requires slightly more caching than the baseline. Our simplification approach may still be ill-timed, since the resimplifications take an additional toll in terms of running time.
@@ 8 Conclusions
We contributed a rigorous provable theory of adaptive multilevel simplification that accelerates solution of belief-dependent fully continuous POMDP. Our theory always identifies the same optimal action or policy as the unsimplified analog. Our theoretical approach receives as input adaptive bounds over the belief dependent reward. Using the suggested theory and these bounds we formulated three algorithms, considering a given belief tree and an anytime MCTS setting. We also contributed a specific simplification considering nonparametric beliefs represented by weighted particles, and derived novel bounds over a differential entropy estimator, which only utilize a subset of the particles and are computationally cheaper than the latter. Our experiments demonstrate that our algorithms are paramount in terms of computation time while guaranteed to have the same performance as the baselines. In the setting of given belief tree we achieved a speedup up to $70\%$. In an anytime MCTS setting, our algorithm enjoyed the speedup of $20\%$.
@@ 10.1 Proof for Theorem 1
To shorten the notations we prove the theorem for value function under arbitrary policy. Note that by substituting the policy $\pi_{(\ell)+}$ by $\{\pi_{\ell}(b_{\ell}),\pi^{*}_{(\ell+1)+}\}$ where $a_{\ell}=\pi_{\ell}(b_{\ell})$ we always can obtain action-value function. Without loosing generality assume the resimplifciation hits an arbitrary belief action node. The new upper bound will be
@@ 10.2 Proof of Lemma 1
Assume in contradiction that the resimplifciation strategy does not promote any reward level and $G(ha)>0$. This means that $\nicefrac{{G(ha)}}{{d}}>0$ and for all reward bounds the inequality $\gamma^{d-d^{\prime}}\cdot(\overline{\rho}-\underline{\rho})<\frac{1}{d}G(ha)$. This is not possible since $\nicefrac{{G(ha)}}{{d}}$ is the mean gap with respect to simulations of MCT and the depth of the belief tree, multiplied by the appropriate discount factor, over all the nodes that are the descendants to $ha$ See equation (32). $\blacksquare$

### v2

@@ 8.2 Entropy Estimators and Bounds Study
In this section, we experiment with a passive case of the continuous 2D Light Dark problem from Section 8.1.1. Our goal is to study the various entropy estimators and our derived bounds from Section 6 over the estimator developed in Boers et al. (2010). In this study, we manually supply the robot with an action sequence to conduct. This results in a single lace of the beliefs corresponding to observations that the robot actually obtained by executing a given externally action sequence. We also provide some attempt in this section to compare estimated reward with the exact analytical counterpart.
@@
The results for the first and second scenarios are provided in Figs. 10 and 11, respectively. For both scenarios we observe that the bounds become tighter as the number of particles of simplified belief $n^{s}_{x}$ increases. We also witness that all estimators vary but the overall trend is similar, putting aside the discrete entropy over the weights. The discrete entropy over the weights fails to adequately represent the uncertainty of the belief. This is an anticipated result. Let us proceed to the planning experiments.
@@ 8.3.1 Acceleration measures
$\frac{\sum_{i}\Big{(}n_{i,x}^{2}-n_{i,x}^{s}n_{i,x}\Big{)}}{\sum_{i}n_{i,x}^{2}}\cdot 100,$ (64)
where the summation is over the future posterior beliefs in all the belief trees in a number of a consecutive planning sessions in particular scenario. Eq. (64) measures relative speedup without time spent on resimplifications. It is calculated at the end of several consecutive planning sessions. To calculate speedup according to (64) one shall pick up the final number of particles of simplified belief $n^{s}_{i,x}$ used for the simplified reward for each belief node $i$, sum over all the nodes of the belief trees (given or constructed on the fly) from planning sessions, make a calculation portrayed by (64). Importantly, acceleration measure (64) assumes that time of evaluating the motion and observation models do not vary from one evaluation to another. If the number of belief particles is not not depending on the belief ($n_{i,x}=n_{x}$) we can further simplify the (64) to
$\frac{\sum_{i}\Big{(}n_{x}-n_{i,x}^{s}\Big{)}}{\sum_{i}n_{x}}\cdot 100.$ (65)
@@
If the quantities (64) and (66) are identical we can conclude that there will be no overhead from resimplifications and adapting the bounds. Note also that in the first place it is not clear how many particles $n_{x}$ for belief representation to take. The number of particles $n_{x}$ shall be as large as possible due to fact that we do not know when the belief represented by weighted particles will converge to the corresponding theoretical belief.
@@ 8.3.3 Results for 2D Continuous Target Tracking in the Setting of a Given Belief Tree
Our action space is $\mathcal{A}=\{\rightarrow,\nearrow,\uparrow,\nwarrow,\leftarrow,\swarrow,\downarrow,\searrow,\mathrm{Null}\}$, where action $\mathrm{Null}$ means that agent doesn’t take any action. In this problem we selected the parameters to be $d_{\text{min}}=0.0001$, $\Sigma_{T}=I\cdot\sigma_{T}$ where $\sigma_{T}=0.1$ and $\sigma_{\mathcal{O}}=0.1$, $\gamma=0.95$.
@@ 8.3.4 Experiments with MCTS
In an anytime setting of MCTS we focus on the $2D$-continuous light dark problem from Section 8.1.1. We place a single “light beacon” in the continuous world. Here we changed the reward. The agent’s goal is to get to location $(0,0)$ and execute the terminal action - $\mathrm{Null}$. Executing it within a radius of $0.5$ from $(0,0)$ will give the agent a reward of $200$, and executing it outside the radius will yield a negative reward of $-200$. For all other actions the multi-objective reward function is $\rho(b,a,z,b^{\prime})=-\underset{x\sim b^{\prime}}{\mathbb{E}}[\|x\|_{2}]-\hat{\mathcal{H}}(b,a,z,b^{\prime})$. The agent can move in eight evenly spread directions $\mathcal{A}=\{\rightarrow,\nearrow,\uparrow,\nwarrow,\leftarrow,\swarrow,\downarrow,\searrow,\mathrm{Null}\}$. Motion and observation models, and the initial belief are $\mathbb{P}_{T}(\cdot|x_{,}a)=\mathcal{N}(x+a,\Sigma_{T})$, $\mathbb{P}_{Z}(z|x)=\mathcal{N}(x,\min\{1,\parallel x-x^{b}\parallel^{2}_{2}\}\cdot\Sigma_{\mathcal{O}})$, $b_{0}=\mathcal{N}(x_{0},\Sigma_{0})$ respectively. $x^{b}$ is the 2D location of the beacon and all covariance matrices are diagonal (i.e. $\Sigma=I\cdot\sigma^{2}$).
@@
We selected the following parameters $x_{0}=\begin{pmatrix}-5.5\\ 0.0\end{pmatrix},\Sigma_{0}=\begin{pmatrix}0.2&0.0\\ 0.0&0.2\end{pmatrix},\sigma_{T}=\sigma_{\mathcal{O}}=0.075$. We experiment with $10$ different configurations (rows of Table 7) that differ in $n_{x}$ (number of particles), $L$ (MCTS simulation depth), and ##\##iter (number of MCTS simulation iterations per planning session). Each scenario comprises $10$ planning sessions, i.e. the agent performs up to $10$ planning action-executing iterations. The scenario stops if the best action determined in planning is $\mathrm{Null}$. We repeat each experiment 25 times. In each such repetition we run PFT-DPW and SITH-PFT with the same seed and calculate the relative time speedup in percentage according to (66) where $t_{\mathrm{PFT-DPW}}$ and $t_{\mathrm{SITH-PFT}}$ are running times of a baseline and our methods respectively.
In all different configurations, we obtained significant time speedup of approximately $20\%$ while achieving the exact same solution compared to PFT. In Table 7 we report the mean and standard error of (66) as well as maximum and minimum value. Remarkably, we observe that we never slowdown the PFT-DPW with SITH-PFT. We also present total running times of $25$ repetitions of at most $10$ (the simulation stops if best identified action is $\mathrm{Null}$) planning sessions in Table 8. Note that we divided the total planning time by the number of planning sessions in each repetition.
@@
8.3.5 Localization with Collision Avoidance Solved by MCTS
In this section, we investigate the application of three algorithms, IPFT (Fischer and Tas, 2020), PFT-DPW (Sunberg and Kochenderfer, 2018) and our SITH-PFT encapsulated by Alg. 7. The algorithmic implementation of IPFT boils down to making more simulations inside IPFT with substantially less number of belief particles subsampled from root belief.
Further, we discuss the quality and speed of IPFT. Representation of the belief with a tiny amount of particles induces larger error in differential entropy estimation and other parts of the reward function such as, for example, soft safety reward component in (63). The authors of (Fischer and Tas, 2020) claim that IPFT averages differential entropies calculated from tiny subsets subsampled from the particle belief. However, observing the SIMULATE routine (similar to our in Alg. 7) in (Fischer and Tas, 2020), we see that in practice this average is obtained through more simulations, starting from a new subsample from the root belief, with less number of particles, thereby averaging entropies calculated from different beliefs with less number of particles, but same history of actions and observations. The parameter $K$ in (Fischer and Tas, 2020) in practice is the visitation count $N(b)$ of each belief in the belief tree. There is no direct control of this parameter. In other words, to make a proper comparison we shall increase the number of SIMULATE calls inside IPFT by a factor $K=\nicefrac{{n_{x}}}{{m}}$ where $m$ is the size of the subsample from a belief represented by $n_{x}$ particles. In such a way in both belief trees, built by IPFT and PFT-DPW, there are the same number of total particles. This is in contrast to using the same number of calls to SIMULATE in both trees. If the number of calls to SIMULATE is the same the number of particles in the tree built by IPFT will be much smaller than in the tree built by PFT-DPW. Do note that we cannot assure that the same $K$ will be for each future history due to the exploratory nature of MCTS.
The speed of IPFT is linked with the rollout policy of MCTS. As we mentioned above, when the belief is represented by particles we know that asymptotically when the number of particles tends to infinity this representation converges to the theoretical belief for any given belief (Crisan and Doucet, 2002). Therefore, we shall take as many particles as possible for the belief representation. Given that the size of subsample $m$ in IPFT does not change, this will increase the parameter $K$ and therefore slowdown IPFT. Because when the new belief node is expanded in the belief tree there is always a rollout initiated, a more complex rollout policy will slowdown IPFT more, yet, this is ultimately the question of how big the parameter $K$ is.
Figure 17: The plot shows $15$ differently colored robot trajectories. Each such trajectory comprises ten time steps. In each such step the robot performs re-planning and executes the best action selected by an appropriate BSP algorithm. The color of each trajectory matches planning with the same seed in each plot. The canvas color here is $\sigma_{\mathcal{O}}=\sigma_{T}=0.03$ as in equations (60) and (61) respectively. The parameters are $n_{x}=300$, $m=20$, number of calls to SIMULATE of IPFT is $4500$, the number of calls to SIMULATE of PFT-DPW and SITH-PFT is $300$. In such a setting the constructed belief trees by these methods have the same number of total samples (see Section 8.3.5 for details). (a) Safe IPFT. (b) Safe SITH-PFT (Alg 7), (c) Safe PFT-DPW.
Figure 18: This plot is associated with Fig. 17. Each color matches the corresponding trajectory in Fig. 17. The parameters are $n_{x}=300$, $m=20$, $K=\frac{300}{20}=15$ number of calls to SIMULATE of IPFT is $4500$, the number of calls to SIMULATE of PFT-DPW and SITH-PFT is $300$. In such a setting the constructed belief trees by these methods have the same number of total samples (see Section 8.3.5 for details). (a) Cumulative information reward as in (63) in the execution of the trajectory. Here the SITH-PFT curve and the PFT-DPW curve overlap. This is because the rewards are identical since the same best action is calculated by SITH-PFT and PFT-DPW; (b) Average planning times of $10$ planning sessions in each trial.
As we observe in Fig. 17, IPFT is less accurate compared to PFT-DPW and SITH-PFT in spite of a much larger number of calls to SIMULATE routine compared to PFT-DPW and SITH-PFT. Clearly, better localization is closer to the beacons. In Fig. 17 we see that more trajectories went to completely different from beacons directions as opposed to Fig. 17 and Fig. 17 displaying identical results. From Fig. 18 we conclude that in $10$ from $15$ trials the information reward obtained in execution of the optimal action returned by IPFT was inferior to the corresponding reward obtained by SITH-PFT and PFT-DPW. From Fig. 18 we see that IPFT is slowest from the three algorithms while SITH-PFT (Alg. 7) is the fastest in all trials.
@@ 8.4 Discussion
Although the speedup was significant and steady for all simulations, we did not observe growth in speed-up with growth of number of belief particles in any simulation. This can be explained by the fact that increasing number of particles of the belief ($n_{x}$) changes also the bounds because the parameter $n_{x}$ is present in the bounds as well. The limitation of our approach is that it leans on converging bounds, which are not trivial to derive and specific for a particular reward function. In addition, it requires slightly more caching than the baseline. Our simplification approach may still be ill-timed, since the resimplifications take an additional toll in terms of running time.
@@ 9 Conclusions
We contributed a rigorous provable theory of adaptive multilevel simplification that accelerates the solution of belief-dependent fully continuous POMDP. Our theory always identifies the same optimal action or policy as the unsimplified analog. Our theoretical approach receives as input adaptive bounds over the belief-dependent reward. Using the suggested theory and any bounds satisfying stated conditions we formulated three algorithms, considering a given belief tree and an anytime MCTS setting. We also contributed a specific simplification for nonparametric beliefs represented by weighted particles and derived novel bounds over a differential entropy estimator. These bounds are computationally cheaper than the latter. Our experiments demonstrate that our algorithms are paramount in terms of computation time while guaranteed to have the same performance as the baselines. In the setting of the given belief tree, we achieved a speedup up to $70\%$. In an anytime MCTS setting, our algorithm enjoyed the speedup of $20\%$.
@@ 11.1 Proof for Theorem 1
To shorten the notations we prove the theorem for value function under arbitrary policy. Note that by substituting the policy $\pi_{(\ell)+}$ by $\{\pi_{\ell}(b_{\ell}),\pi^{*}_{(\ell+1)+}\}$ where $a_{\ell}=\pi_{\ell}(b_{\ell})$ we always can obtain action-value function. Without loosing generality assume the resimplification hits an arbitrary belief action node. The new upper bound will be
@@ 11.2 Proof of Lemma 1
Assume in contradiction that the resimplification strategy does not promote any reward level and $G(ha)>0$. This means that $\nicefrac{{G(ha)}}{{d}}>0$ and for all reward bounds the inequality $\gamma^{d-d^{\prime}}\cdot(\overline{\rho}-\underline{\rho})<\frac{1}{d}G(ha)$. This is not possible since $\nicefrac{{G(ha)}}{{d}}$ is the mean gap with respect to simulations of MCT and the depth of the belief tree, multiplied by the appropriate discount factor, over all the nodes that are the descendants to $ha$ See equation (33). $\blacksquare$